OpenStudy (anonymous):

I need help I forgot how to do (b+6)(b+4) multiply

OpenStudy (shamim):

b(b+4)=?

OpenStudy (shamim):

if u feel difficulty to calculate lz tell me b(b+4)=b^2+4b

OpenStudy (shamim):

now tell me 6(b+4)=?